A 6 inch standard horizontal pipe delivers water at a velocity of 10ft/s to a pond from a river. The length of the pipeline is 1000 ft, if the water temperature is 80F, calculate the theoretical horsepower of the pump.
Answer #1

The total power needed to pump a fluid depende opon the volumetric flow rate of pt fluid, prictional losses, eleration and spGiven data : Nominal pipe size =6"

Temperature of fluid = 80F

Velocity = 10ft/s

Nominal Size (in.) Kw 6.0 7.0 8.0 9.0 Head Loss (ft./100 ft.) Velocity (ft./sec.) 0.04 0.03 0.02 0.01 ZIITT 0.15/03 | 0.2 0.4Frictional losses(FL)=4.5 × 10 = 45 ft

Flow rate (Q) = 950 GPM

Temperature Pressure Density SG at reference temperature [°C] [g/cm3] 4 °C 15 °C 15.6°C 0.1 15 20 25 30 35 40 45 50 55 1 atmspecific gravity at 80F = 0.9718

Theoretical pump horsepower= Q×SG×FL÷ 3960

= 950×45×0.9718÷3960

=10.49 HP